Gucchi:

Jacob makes the following statement: "I have a 60% chance of being selected as the captain of the baseball team and a 20% chance of being elected student body president." What is the probability that Jacob will be selected as the captain of the baseball team and will be elected student body president? 80% 40% 12% 8%

Gucchi:

I'm kinda confused, is this B or C?

AZ:

So when you have the word AND, you have to multiply when you have the word OR, you would add so Probability of being captain = 60% Probability of being president = 20% probability of being captain AND president = 60% * 20%

Gucchi:

So I would do 60 x 20?

AZ:

careful, remember that it's 60 PERCENT and percent means over 100 60% = 60/100 = 0.6

AZ:

same thing for 20%

Gucchi:

Okay

Gucchi:

So 0.6 x 0.2?

snowflake0531:

Yep, adn then multiply by 100

AZ:

yes, and you'll get a decimal and then to make it into a percentage, you'll multiply by 100

Gucchi:

I multipled 0.6 x 0.2 and got 0.12, then i multiplied by 100 to get 12

AZ:

You got it!!

Gucchi:

Okay, I get it. Thank you

AZ:

To be clear, we're multiplying by 100% remember how percent is over 100 100% would be 100/100 so we're basically multiplying by 1 and so when you multiply by 100% 0.12 * 100% = 12%

AZ:

so essentially, you just multiply by 100 but you keep that percent symbol still with it

Gucchi:

Yeah, i will remmeber that.

Gucchi:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ so essentially, you just multiply by 100 but you keep that percent symbol still with it \(\color{#0cbb34}{\text{End of Quote}}\) Okay.